OpenStudy (anonymous):

-3(x-2) is less than or equal to 1/3 Solve the inequality Verify the solution describing the solution to the inequality Verify your solution to the inequality using two elements of the solution set

OpenStudy (texaschic101):

-3(x - 2) <= 1/3 -- distribute through the parenthesis -3x + 6 <= 1/3 --- subtract 6 from both sides -3x <= 1/3 - 6 -3x <= 1/3 - 18/3 -3x <= - 17/3 -- divide both sides by -3, and change the sign x >= (-17/3) / (-3) x >= -17/3 * -1/3 x >= 17/9 or 1 8/9 (thats 1 and 8/9)